# Relative Pose Estimation for Nonholonomic Robot Formation with UWB-IO Measurements

### v2

@@ VI Conclusion
$\dot{\textbf{p}}^{\mathcal{O}_{0}}_{i}(t)&=\begin{bmatrix}\cos\phi_{\Sigma_{i}}^{\mathcal{O}_{0}}(t),&0\\ \sin\phi_{\Sigma_{i}}^{\mathcal{O}_{0}}(t),&0\\ 0,&1\end{bmatrix}\begin{bmatrix}v_{i,h}\\ v_{i,z}\end{bmatrix}\\ \dot{\phi}_{\Sigma_{i}}^{\mathcal{O}_{0}}(t)&=w_{i}.\end{aligned}\right.$ (35)
@@
-D Complete observability analysis of relative localization
@@
$t_{k}$pbox10cmUWB & IO measurement time instant for robot $i$.
$\mathcal{O}_{i}$pbox10cmThe odometry frame of robot $i$.
$\Sigma_{i}$pbox10cmReal-time IMU body frame of robot $i$.
$\textbf{p}^{\mathcal{O}_{i}}_{i}$pbox10cmDisplacement of robot $i$ in frame $\mathcal{O}_{i}$ from $t_{0}$ to $t_{k}$.
$\phi_{\Sigma_{i}}^{\mathcal{O}_{i}}(t)$pbox10cmAngular displacement of robot $i$ in frame $\mathcal{O}_{i}$ from $t_{0}$ to $t_{k}$.
$\textbf{u}^{\mathcal{O}_{i}}_{i}(t_{k})$pbox10cmDisplacement of robot $i$ in frame $\mathcal{O}_{i}$ from $t_{k}$ to $t_{k+1}$
$d_{ij}(t_{k})$pbox10cmDistance between robots $i$ and $j$.
$y_{ij}$pbox10cmAuxiliary signal which records innovation.
$\hat{\textbf{p}}^{\mathcal{O}_{i}}_{ij,0}$pbox10cm Estimate of initial relative position between robot $i$ and its
neighbor robot $j$ in frame $\mathcal{O}_{i}$.
$\hat{\textbf{R}}^{\mathcal{O}_{i}}_{\mathcal{O}_{j}}$pbox10cm Estimate of initial rotation matrix between frame $\mathcal{O}_{i}$ and $\mathcal{O}_{j}$.
$\hat{\textbf{q}}^{\mathcal{O}_{i}}_{i0}$pbox10cm Estimate of initial relative position between robot $i$ and leader
robot 00 in frame $\mathcal{O}_{i}$.
$\hat{\textbf{Q}}^{\mathcal{O}_{i}}_{\mathcal{O}_{0}}$pbox10cm Estimate of rotation matrix between Frame $\mathcal{O}_{i}$ and $\mathcal{O}_{0}$.
$v_{i}$pbox10cmVelocity control command for robot $i$.
[…]
@@
Appendix A Technical details
A-A Notation list of the proposed method
@@
A-B Pseudocode of the proposed method
The pseudocode of the proposed data-driven localization and formation control method is shown in 1.
@@ A-B Pseudocode of the proposed method
1 while $k+1$-th measurement time instant do
@@ 16             Calculate control command with (24
1 Judgment data queue $Q_{ij,\mathrm{judge}}$leftarrow empty set $\phi$;
@@
3       Current data triplet $\mathcal{D}_{ij}[k]$leftarrow UWB & IO measurement;
4       Outlier vote number $n_{\mathrm{outlier}}$leftarrow 0;
@@
7                   $n_{\mathrm{outlier}}$leftarrow $n_{\mathrm{outlier}}+1$;
8
@@ 7                   $n_{\mathrm{outlier}}$leftarrow $n_{\mathrm{outlier}}+1$;
10      Data queue $n_{\mathrm{judge}}$leftarrow data triplets number in $Q_{ij,\mathrm{judge}}$;
11       Outlier detection flag initialize $f_{ij,\mathrm{outlier}}$leftarrow 1;
@@
13             $f_{ij,\mathrm{outlier}}$leftarrow 0
@@ 15             Update $Q_{ij,\mathrm{judge}}$ with the current Measurement data triplet $\mathcal{D}_{ij}[k]$
Appendix B Supplemental simulation and experiment results
B-A Discussion on parameter setting
@@
Figure 19: Velocity and angle velocity for each robot in the swarm.
@@
B-B Motion trajectory of comparative simulation
@@
B-C Formation control for large scale swarm
@@
B-D Noise robustness for large scale swarm
@@
B-E Sensor highly inaccurate case
@@
B-F Formation control with different initial state

### v4

@@ A-C Proof of Theorem 3
$\left\{\begin{aligned} \dot{\textbf{p}}^{\mathcal{O}_{0}}_{i}(t)&=\begin{bmatrix}\cos\phi_{\Sigma_{i}}^{\mathcal{O}_{0}}(t),&0\\ \sin\phi_{\Sigma_{i}}^{\mathcal{O}_{0}}(t),&0\\ 0,&1\end{bmatrix}\begin{bmatrix}v_{i,h}\\ v_{i,z}\end{bmatrix}\\ \dot{\phi}_{\Sigma_{i}}^{\mathcal{O}_{0}}(t)&=w_{i}.\end{aligned}\right$. (35)
@@
A-D Complete observability analysis of relative localization
@@
$t_{k}$ UWB & IO measurement time instant for robot $i$.
$\mathcal{O}_{i}$ The odometry frame of robot $i$.
$\Sigma_{i}$ Real-time IMU body frame of robot $i$.
$\textbf{p}^{\mathcal{O}_{i}}_{i}$ Displacement of robot $i$ in frame $\mathcal{O}_{i}$ from $t_{0}$ to $t_{k}$.
$\phi_{\Sigma_{i}}^{\mathcal{O}_{i}}(t)$ Angular displacement of robot $i$ in frame $\mathcal{O}_{i}$ from $t_{0}$ to $t_{k}$.
$\textbf{u}^{\mathcal{O}_{i}}_{i}(t_{k})$ Displacement of robot $i$ in frame $\mathcal{O}_{i}$ from $t_{k}$ to $t_{k+1}$
$d_{ij}(t_{k})$ Distance between robots $i$ and $j$.
$y_{ij}$ Auxiliary signal which records innovation.
$\hat{\textbf{p}}^{\mathcal{O}_{i}}_{ij,0}$ Estimate of initial relative position between robot $i$ and its neighbor robot $j$ in frame $\mathcal{O}_{i}$.
$\hat{\textbf{R}}^{\mathcal{O}_{i}}_{\mathcal{O}_{j}}$ Estimate of initial rotation matrix between frame $\mathcal{O}_{i}$ and $\mathcal{O}_{j}$.
$\hat{\textbf{q}}^{\mathcal{O}_{i}}_{i0}$ Estimate of initial relative position between robot $i$ and leader robot 0 in frame $\mathcal{O}_{i}$.
$\hat{\textbf{Q}}^{\mathcal{O}_{i}}_{\mathcal{O}_{0}}$ Estimate of rotation matrix between Frame $\mathcal{O}_{i}$ and $\mathcal{O}_{0}$.
$v_{i}$ Velocity control command for robot $i$.
When robot $i$ and its neighbor robot $j$ remain relatively stationary all the time, i.e. $v_{i}^{\mathcal{O}_{i}}\equiv v_{j}^{\mathcal{O}_{j}}$, at least the z-axis relative position is unobservable. One can see that if robot $i$ and its neighbor robot $j$ remain relatively stationary all the time, then $u^{\mathcal{O}_{i}}_{i,z}(t_{k})=u^{\mathcal{O}_{j}}_{j,z}(t_{k})$, then $S_{ij}=\begin{bmatrix}\textbf{{*}}_{2\times 2},&0,&\textbf{*{0}}_{2\times 4}\\ \textbf{0}_{1\times 2},&0,&\textbf{0}_{1\times 4}\\ \textbf{*}_{4\times 2},&0,&\textbf{0}_{4\times 4}\end{bmatrix}$. The data matrix block corresponding to the z-axis relative position is always 0, indicating that it is unobservable.
@@
Appendix B Technical details
B-A Notation list of the proposed method
@@
B-B Pseudocode of the proposed method
The pseudocode of the proposed data-driven localization and formation control method is shown in Algorithm 1.
@@ B-B Pseudocode of the proposed method
1
21exwhile $k+1$-th measurement time instant do
@@ 17     Calculate control command with (24
Input: Robot ID $i$, Neighbor set $\mathcal{N}_{i}$, Neighbor hop $h_{j}$, Leader information estimated by neighbor $\hat{\sigma}_{j}$
Output: Estimatin of leader information $\hat{\sigma}_{i}$
1
[…]
3 while $k$-th measurement time instant do
4  if ID $i$ == leader ID 0 then
5     $\hat{\sigma}_{i}$ $\leftarrow$ $\sigma_{i}$;
6     $h_{i}$ $\leftarrow$ 0;
7
8 else
[…]
10     $h_{i}$ $\leftarrow$ $h_{j}+1$;
11     $\hat{\sigma}_{i}$ $\leftarrow$ $\sigma_{j}$;
12
13
Algorithm 2 Multi hop propagation algorithm for leader information
In engineering, in addition to the multi hop propagation method, consensus estimation can also be used to achieve consistent estimation of the leader odometer information for each robot. Specifically, the leader odometer information is used as a reference signal, and a consensus observer is designed by constructing a consensus error to achieve asymptotic estimation of the time-varying signal of the leader odometer information.
@@ 11     $\hat{\sigma}_{i}$ $\leftarrow$ $\sigma_{j}$;
1
21exJudgment data queue $Q_{ij,\mathrm{judge}}$ $\leftarrow$ empty set $\phi$;
@@
4  Current data triplet $\mathcal{D}_{ij}[k]$ $\leftarrow$ UWB & IO measurement;
5  Outlier vote number $n_{\mathrm{outlier}}$ $\leftarrow$ 0;
@@
8        $n_{\mathrm{outlier}}$ $\leftarrow$ $n_{\mathrm{outlier}}+1$;
@@
10
11 Data queue $n_{\mathrm{judge}}$ $\leftarrow$ data triplets number in $Q_{ij,\mathrm{judge}}$;
12  Outlier detection flag initialize $f_{ij,\mathrm{outlier}}$ $\leftarrow$ 1;
@@
14     $f_{ij,\mathrm{outlier}}$ $\leftarrow$ 0
@@ 16     Update $Q_{ij,\mathrm{judge}}$ with the current Measurement data triplet $\mathcal{D}_{ij}[k]$
B-C Computation time and communication bandwidth
TABLE II: Performance Comparison on Different Hardware Platforms
Metric Laptop (X86-64) Jetson Orin (ARM)
8 Cores, 2.7 GHz 8 Cores, 2.0 GHz
Minimum program
loop time (s)
0.000144 0.000459
Maximum program
loop time (s)
0.002904 0.008829
Average program
loop time (s)
0.000160 0.000481
Standard deviation
of program loop time (s)
0.000035 0.000089
A discussion on the algorithm’s communication bandwidth requirements and computation time is presented in this section. Taking robot $i$ as an example, in each algorithm iteration, the variable information of neighbor $j$ needs to be used. According to the methods mentioned in the paper, they roughly include: inter-robot distance $d_{ij}(t_{k})$ (1 float); odometer displacement of robot $j$ $\textbf{p}^{\mathcal{O}_{j}}_{j}(t_{k})$ (3 float); odometer angle displacement of robot $j$ $\phi_{\Sigma_{j}}^{\mathcal{O}_{j}}(t_{k})$ (1 float); hop count of robot $j$ (1 float); leader estimator $\hat{\textbf{q}}^{\mathcal{O}_{j}}_{j0}$ (3 float) and $\hat{\textbf{Q}}^{\mathcal{O}_{j}}_{\mathcal{O}_{0}}$ (2 float) of robot $j$; leader odometer information for multi hop propagation $v_{0,h}(t_{k})$ (1 float); $v_{0,z}(t_{k})$ (1 float); $w_{0}(t_{k})$ (1 float); $\textbf{p}^{\mathcal{O}_{0}}_{0}(t_{k})$ (3 float); $\phi_{\Sigma_{0}}^{\mathcal{O}_{0}}(t_{k})$ (1 float). At time instant $t_{k}$, for robot $i$ and robot $j$, communication and interaction of $18\times 4=64$ Byte of information is required. Assuming that robot $i$ has 10 neighbors and the algorithm iteration frame rate is 20 frames per second (i.e., $\Delta t=0.05$s, consistent with our simulation and experiment), the communication bandwidth required by robot i is $64\times 10\times 20/1024=12.5$kb/s. According to the calculation results, it can be known that increasing the iteration frequency of the algorithm will require higher communication bandwidth. However, it should be pointed out that the algorithm execution frequency of 20 frames is generally sufficient to meet the localization and control requirements. In practical use, the algorithm calculation frame rate can be appropriately increased under the condition of meeting the communication bandwidth. Besides, even if the swarm size is large, due to the distributed algorithm, it only needs to communicate and interact with some neighboring robots. The existing UWB module can fully meet this bandwidth. In addition, to verify the computational speed of the algorithm on the onboard measurement platform (jetson orin), we ran the same simulation experiment code simultaneously on a laptop and jetson platform, recording the time required for each iteration of the algorithm (including data collection, data filtering, auxiliary variable calculation, relative localization algorithm iteration, cooperative localization calculation, etc.). The specific results are shown in the table II.
Appendix C Supplemental simulation and experiment results
C-A Discussion on parameter setting
@@
Figure 19: Speed and angle velocity for each robot in the swarm.
@@
C-B Motion trajectory of comparative simulation
@@
C-C Formation control for large scale swarm
@@
Figure 20: Pose estimation error and formation control error for each robot in the swarm with multi hop propagation of leader odometer measurement information. A: Estimation error. B: Formation control error.
Figure 21: Speed and angle velocity for each robot with multi-hop propagation of leader odometer measurement information.
In order to verify the method of multi hop propagation of leader odometer measurement information, we conducted simulation experiments in a random topology consisting of 50 robots. The experimental results are shown in Fig. 20 and Fig. 21, which is basically consistent with directly using leader information (Fig. 18 and Fig. 19). This is because in engineering, a neighbor communication frame rate of 20 frames is sufficient for each robot to observe leader odometer information in a very short time (robot’s own layers $\times$ 0.05 seconds).
C-D Noise robustness for large scale swarm
@@
C-E Sensor highly inaccurate case
@@
C-F Formation control with dynamic topology
For the dynamic topology problem, we conducted simulation experiments under dynamic switching topology conditions. Specifically, 50 robots had random initial poses and switched measurement topologies at 400 seconds and 800 seconds, respectively. The experimental results are shown in Fig. 24 and Fig. 25. It can be seen that after switching topologies, the robots re positioned their neighbors and carried out cooperative localization and distributed formation of leader robots based on this. The experimental results show that the algorithm converges quickly after topology dynamic switching.
Figure 24: Pose estimation error and formation control error for each robot in the swarm under switched measurement topologies. A: Estimation error. B: Formation control error.
Figure 25: Velocity and angle velocity for each robot under switched measurement topologies.
C-G Formation control with different initial state